### v2

# Photon propagation in a charged Bose-Einstein condensate model

Abstract
We consider the propagation of photons in a model of a charged scalar Bose-Einstein (BE) condensate. We determine the dispersion relations of the collective modes, as well as the photon polarization tensor and the dielectric constant in the model. Two modes correspond to the transverse photon polarizations, with dispersion relations of the usual form for transverse photons in a plasma. The other two modes, denoted as the (±)plus-or-minus($\pm$)( ± ) modes, are combinations of the longitudinal photon and the massive scalar field. Their dispersion relations behave very differently as functions of momentum. The (+)(+)( + ) mode dispersion relation increases steadily and remains greater than the momentum as the momentum increases. The dispersion relation of the ()(-)( - ) mode decreases in a given momentum range, with the group velocity being negative in that range, while in another range it increases steadily but remains smaller than the momentum, akin to the situation in a medium with an index of refraction greater than 1. We consider the non-relativistic limit of the (±)plus-or-minus($\pm$)( ± ) dispersion relations and discuss some aspects of the results. We also determine the wavefunctions of the (±)plus-or-minus($\pm$)( ± ) modes, which are useful to obtain the corrections to the dispersion relations, e.g., imaginary parts due to the damping effects and/or the effects of scattering, due to the interactions with the excitations of the system. The results can be useful in various physical contexts that have been considered in the literature involving the electrodynamics of a charged scalar BE condensate.
1 Introduction and motivation
Many extensions of the standard electroweak theory depend on the presence of additional scalar particles, both neutral and charged, that interact with the standard model particles, in particular with the neutrinos and photons. Such interactions typically have implications in the context of nuclear, astrophysics, and cosmological plasmas that may contain a background of the scalar particles. As is well known, the intrinsic properties of neutrinos and photons propagating in such a medium, such as the dispersion relations and/or electromagnetic properties, are modified in ways that lead to remarkable and observable effects.
In previous works we have carried out a systematic calculation of the neutrino dispersion relation, including the damping and decoherence effects in models in which the neutrinos interact with a scalar ($\phi$) and fermion ($f$) via a coupling of the form $\bar{f}_{R}\nu_{L}\phi$, or just with neutrinos themselves $\bar{\nu}^{c}_{R}\nu_{L}\phi$ (see Ref. [1] and references therein). Couplings of those forms produce additional contributions to the neutrino effective potential when the neutrino propagates in a background of $\phi$ and $f$ particles, or a pure neutrino background, which can effects in various neutrino physics contexts. More recently, we have considered the problem of determining the corresponding quantities (e.g, effective potential and/or dispersion relation and damping) of a fermion that propagates in a thermal background that contains a scalar Bose-Einstein (BE) condensate[2]. The method can be applied to the case of a neutrino propagating in such a background, and the problem of fermions propagating in such backgrounds can be relevant in other applications in astrophysical, heavy-ion collisions and dark-matter contexts[3, 4, 5, 6, 7, 8, 9].
Our motivation in the present work is to consider, in a similar vein, the propagation of photons in a thermal background that contains a BE condensate of a charged scalar. The problem of the propagation of photons in a medium is a classic subject which in the language of plasma physics can be treated within the formalism of the kinetic Boltzmann equation, and in other contexts and/or more generally using the framework of thermal field theory (TFT). Various aspects and applications of the problem of photon propagation in a BE condensate of the charged scalars have been studied in previous recent works, in various contexts and using different approaches, for example the electrodynamics of fermions and charged scalar bosons at low temperature[10], heavy-ion collisions[11], exciton-polaritons[12], the BE condensation of photons[13, 14] and the Cerenkov effect[15], and the reduction of the photon velocity[16], in atomic gases.
Our goal in the present work is to extend the technique used in Ref. [2] to the case in which the scalar is electromagnetically charged and study the propagation of photons in the BE condensate background of such scalars. The field theoretical method we used in Ref. [2] to treat the BE condensate, which we use here, has been discussed by various authors[17, 18, 19, 20]. As we will see, this method allows us to determine succinctly the dispersion relations of the propagating modes after the symmetry breaking associated with the BE condensation in the model, as well as the corresponding polarization tensor and dielectric constant. The method also allows us to determine in a very efficient way the mixing parameters between the longitudinal photon (plasmon) and the physical scalar field that make up the propagating modes involving the plasmon. These are relevant for the calculation of thermal corrections to the dispersion relations (e.g., imaginary parts due to damping effects) and the other physical quantities due to interactions with the thermal excitations of the BE condensate. The results described above are obtained in a general way, within the context of the model, but not tied to any specific application. In this way they can be useful and pave the way for further development in various contexts and applications such as those considered in the references mentioned above. While the principles and techniques used here to treat the charged BE condensate share some common ingredients with those of Ref. [20], our scope, discussion and outlook for applications are very different.
The plan of the presentation describing the present work is as follows. In Section 2 we summarize the essential features and the symmetry breaking mechanism of the scalar BE condensation model discussed in Ref. [2], which serves as the basis and motivation for the extension we present in Section 3 to the charged scalar case. In Section 3 we discuss various qualitative properties of the model that are useful for the interpretation, such as current conservation, symmetry breaking and the spectrum. The propagating modes and the corresponding dispersion relations are determined in Section 4. As we show, two modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. The other two modes, which we denote as the (±)plus-or-minus($\pm$)( ± ) modes, are combinations of the longitudinal photon and the massive scalar field. Various distinguishing aspects of the results obtained are discussed in Section 5. While the dispersion relation of the (+)(+)( + ) mode increases steadily and remains greater than the momentum as the momentum increases, the dispersion relation of the ()(-)( - ) mode reveals some special features. A particularly noticeable one is the fact that it decreases in a given momentum range, and the corresponding group velocity is negative in that range. In another range, it increases steadily for values of the momentum above a given threshold, but such that it remains smaller than the momentum, reminiscent of a medium with index of refraction greater than 1. We also consider the non-relativistic limit of the dispersion relations and discuss some aspects of the results obtained. In particular, under some conditions, the dispersion relation of the ()(-)( - ) mode reduces to the expression for the dispersion relation that has been obtained in non-relativistic models of the BE condensation of charged scalars[21, 22]. In Section 6 we obtain the expression for the photon polarization tensor, or self-energy. On one hand, this provides an alternative way to obtain the dispersion relations. More importantly, it has the virtue that it gives the contribution to the dielectric function of the system due to the BE condensate, which is useful on its own merits, for example for studying also the response of the system to external fields. Finally in Section 7 we determine the mixing parameters of the spinors of the (±)plus-or-minus($\pm$)( ± ) modes. These parameters can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system. In Section 8 we summarize our conclusions and outlook.
2 Scalar BE condensate model
As a preamble to the presentation of the model for the charged scalar coupled to the photon, we first summarize the essential features of the scalar BE condensation model as well as its justification.
In Ref. [2] we used the following model for the $\phi$,
$L_{\phi}=(d^{\mu}\phi)^{\ast}(d_{\mu}\phi)-V_{\phi}\,,$ (2.1)
where
$d_{\mu}\equiv\partial_{\mu}-i\mu u_{\mu}\,,$ (2.2)
and
$V_{\phi}=m^{2}\phi^{\ast}\phi+\lambda(\phi^{\ast}\phi)^{2}\,.$ (2.3)
$\mu$ is a parameter, to be identified with the chemical potential of the $\phi$ and $u^{\mu}$, which is to be identified with the velocity four-vector of the medium, is
$u^{\mu}=(1,\vec{0})\,,$ (2.4)
in the medium’s own rest frame. To simplify the notation, we write
$v_{\mu}=\mu u_{\mu}\,.$ (2.5)
Expanding the $d$ term in Eq. (2.1), the Lagrangian density is
$L_{\phi}=(\partial^{\mu}\phi)^{\ast}(\partial_{\mu}\phi)+i[\phi^{\ast}(v\cdot\partial\phi)-(v\cdot\partial\phi^{\ast})\phi]-U(\phi)\,,$ (2.6)
where
$U\equiv V_{\phi}-\mu^{2}\phi^{\ast}\phi=-(\mu^{2}-m^{2})\phi^{\ast}\phi+\lambda(\phi^{\ast}\phi)^{2}\,.$ (2.7)
If $m^{2}>\mu^{2}$, this $U$ corresponds to a standard massive complex scalar with mass $m^{2}-\mu^{2}$. On the other hand, if $\mu^{2}>m^{2}$, the minimum of the potential is not at $\phi=0$, and therefore $\phi$ develops a non-zero expectation value and the $U(1)$ symmetry is broken.
We assume the second option,
$\mu^{2}>m^{2}\,,$ (2.8)
and proceed accordingly. Namely, we put
$\phi=\frac{1}{\sqrt{2}}\left(\phi_{0}+\phi_{1}+i\phi_{2}\right)\,,$ (2.9)
where
$\langle\phi\rangle\equiv\frac{1}{\sqrt{2}}\phi_{0}\,,$ (2.10)
is chosen to be the minimum of
$U_{0}=-\frac{1}{2}(\mu^{2}-m^{2})\phi^{2}_{0}+\frac{1}{4}\lambda\phi^{4}_{0}\,.$ (2.11)
Thus,
$\phi^{2}_{0}=\frac{\mu^{2}-m^{2}}{\lambda}\,.$ (2.12)
Substituting Eqs. (2.9) and (2.12) in Eq. (2.1) we obtain the Lagrangian density for the scalar excitations $\phi_{1,2}$,
$L_{\phi}=\frac{1}{2}\left[(\partial^{\mu}\phi_{1})^{2}+(\partial^{\mu}\phi_{2})^{2}\right]+\phi_{2}v\cdot\partial\phi_{1}-\phi_{1}v\cdot\partial\phi_{2}-U(\phi)\,,$ (2.13)
where
$U(\phi)=-\frac{1}{2}(\mu^{2}-m^{2})[(\phi_{0}+\phi_{1})^{2}+\phi^{2}_{2}]+\frac{1}{4}\lambda[(\phi_{0}+\phi_{1})^{2}+\phi^{2}_{2}]^{2}\,.$ (2.14)
Using Eq. (2.12), the quadratic part of $U$ involves only $\phi_{1}$,
$U_{2}=\frac{1}{2}m^{2}_{1}\phi^{2}_{1}\,,$ (2.15)
where
$m^{2}_{1}=2(\mu^{2}-m^{2})\,,$ (2.16)
while $\phi_{2}$ does not appear. However, $\phi_{1}$ and $\phi_{2}$ are mixed by the $v^{\mu}$ term. Therefore the next step is to find the propagator matrix of the $\phi_{1,2}$ complex and find the modes that have a definite dispersion relation.
A convenient way to do it is to use matrix notation,
$\hat{\phi}=\left(\begin{array}[]{l}\phi_{1}\\ \phi_{2}\end{array}\right)\,.$ (2.17)
The Lagrangian density, in momentum space, can then be written in the form
$L^{(2)}_{\phi}(k)=\frac{1}{2}\hat{\phi}^{\ast}(k)\Delta^{-1}_{\phi}(k)\hat{\phi}(k)\,,$ (2.18)
where
$\Delta^{-1}_{\phi}(k)=\left(\begin{array}[]{ll}k^{2}-m^{2}_{1}&2iv\cdot k\\ -2iv\cdot k&k^{2}\end{array}\right)\,.$ (2.19)
The classical equations of motion are then
$\Delta^{-1}_{\phi}(k)\hat{\phi}=0\,.$ (2.20)
The dispersion relations of the eigenmodes are given by the solutions of
$D=0\,,$ (2.21)
where $D$ is the determinant of $\Delta^{-1}_{\phi}$,
$D=(\omega^{2}-\kappa^{2})(\omega^{2}-\kappa^{2}-m^{2}_{1})-4\mu^{2}\omega^{2}\,.$ (2.22)
In Eq. (2.22) we have used Eq. (2.5) and introduced the variables $\omega$ and $\kappa$,
$\omega$ $=$ $u\cdot k\,,$
$\kappa$ $=$ $\sqrt{\omega^{2}-k^{2}}\,.$ (2.23)
In the rest frame of the medium
$k^{\mu}=(\omega,\vec{\kappa})\,,$ (2.24)
with $\kappa=|\vec{\kappa}|$. The dispersion relations obtained in this way can be written in the form
$\omega^{2}_{\pm}(\kappa)=\kappa^{2}+\frac{1}{2}m^{2}_{\rho}\pm\sqrt{\frac{1}{4}m^{4}_{\rho}+4\mu^{2}\kappa^{2}}\,,$ (2.25)
where
$m^{2}_{\rho}=m^{2}_{1}+4\mu^{2}\,,$ (2.26)
and they satisfy
$\omega_{+}(0)$ $=$ $m_{\rho}\,,$
$\omega_{-}(0)$ $=$ $0\,.$ (2.27)
The combinations of $\phi_{1,2}$ that have the definite dispersion relations $\omega_{\pm}$ can be obtained from Eq. (2.20). The mode corresponding to $\omega_{-}$ is the realization of the Goldstone mode associated with the breaking of the global $U(1)$ symmetry. However, its dispersion relation is not the usual $\omega=\kappa$, but the one given by $\omega_{-}(\kappa)$ in Eq. (2.25).
As discussed in Ref. [2], the motivation for considering this model is based on the following result: the calculation of the effective potential $V_{\text{eff}}(T,\mu)$ for $\phi$ can be carried out in TFT using $\mu=0$ in the partition (and/or distribution) function, but using the $\mu$-dependent Lagrangian density $L_{\phi}$ defined in Eq. (2.1). Then neglecting the $T$-dependent terms (that is, at zero temperature), $V_{\text{eff}}(0,\mu)$ is simply the $U$ potential given in Eq. (2.7). In other words, calculations with the $\phi$ propagators, $\phi$ Lagrangian density and partition function for $\phi$,
$Z=e^{-\beta H+\alpha Q}\,,$ (2.28)
give the same result as calculating them with the Lagrangian density for $\phi^{\prime}$ obtained by using $\partial_{\mu}\rightarrow D_{\mu}=\partial_{\mu}-i\mu u_{\mu}$ and using
$Z=e^{-\beta H^{\prime}}\,.$ (2.29)
The simple way to obtain this result is the following. The evolution equation for $\phi$ is
$i\partial_{t}\phi=-[\cal H,\phi]\,.$ (2.30)
If now calculate the time derivative of $\phi^{\prime}\equiv e^{i\mu t}\phi$ we get
$\partial_{t}\phi^{\prime}$ $=$ $i\mu\phi^{\prime}+e^{i\mu t}\partial_{t}\phi$ (2.31)
$=$ $i\mu\phi^{\prime}+e^{i\mu t}i[\cal H,\phi]$
$=$ $i\mu\phi^{\prime}+i[\cal H,\phi^{\prime}]$
$=$ $-i[\mu Q,\phi^{\prime}]+i[\cal H,\phi^{\prime}]$
$=$ $i[\cal H-\mu Q,\phi^{\prime}]\,.$
Therefore
$i\partial_{t}\phi^{\prime}=-[\cal H-\mu Q,\phi^{\prime}]\,,$ (2.32)
which is the result stated above, with the identification of $\cal H^{\prime}$. On the other hand, when we express the partition function in terms of the field $\phi^{\prime}$, then
$Z\equiv e^{-\beta\cal H+\alpha Q}=e^{-\beta(\cal H-\mu Q)}=e^{-\beta\cal H^{\prime}}\,.$ (2.33)
That is, in the calculations involving $\phi^{\prime}$ we use the partition function with $\cal H^{\prime}$ and zero chemical potential.
The operator proof we have presented seems to be applicable to any field (e.g., a fermion field), not involving Lagrangian dynamics arguments, and therefore the result shown above holds for any field and not just for a scalar field. On the basis of this result, the Lagrangian formulation that we have presented is justified, which is the most efficient way to do the dynamics of the $\phi^{\prime}$, that is finding the evolution equation to obtain the dispersion relations and other properties (e.g., the propagators) of the eigenmodes.
3 Charged scalar BE condensate model
3.1 Lagrangian
We extend the model by assuming that the $\phi$ field is electrically charged and interacts with electromagnetic field in the usual way. The Lagrangian density is then
$L=-\frac{1}{4}F^{2}+L_{\phi A}\,,$ (3.1)
where $F_{\mu\nu}$ is the electromagnetic field tensor and
$L_{\phi A}=(\hat{D}^{\mu}\phi)^{\ast}(\hat{D}_{\mu}\phi)-V_{\phi}\,,$ (3.2)
with[23]
$\hat{D}_{\mu}=d_{\mu}+iqA_{\mu}=\partial_{\mu}-iv_{\mu}+iqA_{\mu}\,,$ (3.3)
and $V_{\phi}$ given as before in Eq. (2.3). $v^{\mu}$ is defined in Eq. (2.5) and we will sometimes write
$\hat{D}_{\mu}=\partial_{\mu}-i\omega_{\mu}\,,$ (3.4)
where
$\omega_{\mu}=v_{\mu}-qA_{\mu}\,.$ (3.5)
3.2 Spectrum of the model
To determine the spectrum of the model we adopt the unitary gauge, which is the convenient one to use for this purpose. Thus we parametrize $\phi$ in the form
$\phi=\frac{1}{\sqrt{2}}(\phi_{0}+\rho)e^{i\theta/\phi_{0}}\,.$ (3.6)
The field $\theta$ does not appear in $V_{\phi}$, and by a gauge transformation it disappears also from the kinetic term. To be clear that we are employing the unitary gauge we will denote by $V_{\mu}$ the transformed vector potential $V_{\mu}=A_{\mu}+\frac{1}{q\phi_{0}}\partial_{\mu}\theta$. It then follows that
$L_{\phi A}=\frac{1}{2}(\partial\rho)^{2}+\frac{1}{2}(v-qV)^{\mu}(v-qV)_{\mu}(\phi_{0}+\rho)^{2}-V_{\phi}\,,$ (3.7)
or, expanding the various factors,
$L_{\phi A}=\frac{1}{2}(\partial\rho)^{2}+\frac{1}{2}(q\phi_{0})^{2}V^{2}+\frac{1}{2}q^{2}V^{2}\rho^{2}+q^{2}\phi_{0}\rho V^{2}-qv\cdot V(\phi_{0}+\rho)^{2}-U_{\phi}\,,$ (3.8)
with
$U_{\phi}=-\frac{1}{2}(\mu^{2}-m^{2})(\phi_{0}+\rho)^{2}+\frac{1}{4}(\phi_{0}+\rho)^{4}\,.$ (3.9)
Proceeding as in Section 2, we assume
$\mu^{2}>m^{2}\,,$ (3.10)
and choose $\phi_{0}$ such that $U_{\phi}$ has the minimum at $\rho=0$, thus,
$\phi^{2}_{0}=\frac{\mu^{2}-m^{2}}{\lambda}\,.$ (3.11)
From Eq. (3.8), the bilinear part of $L$, including the $F^{2}$ term, is then
$L^{(2)}=-\frac{1}{4}F^{2}+\frac{1}{2}m^{2}_{V}V^{2}+\frac{1}{2}(\partial\rho)^{2}-\frac{1}{2}m^{2}_{1}\rho^{2}-2m_{V}(v\cdot V)\rho\,,$ (3.12)
where
$m_{V}$ $=$ $q\phi_{0}\,,$
$m^{2}_{1}$ $=$ $2(\mu^{2}-m^{2})\,.$ (3.13)
The physical picture that emerges is this: the field $\theta$ becomes the longitudinal component of $V$, and we end up with two fields, $V$ and $\rho$. The longitudinal component of $V$ is mixed with the scalar field $\rho$ (the term $\rho v\cdot V)$. As we show, the consequence is that the propagating modes with definite dispersion relations involve a superposition of the longitudinal component of $V$ and the $\rho$.
The following comment is in order before proceeding. In Eq. (3.8) there is a term linear in $V_{\mu}$ of the form $q\phi^{2}_{0}v\cdot V=(q\mu\phi^{2}_{0})u\cdot V$. This term is a tadpole, meaning that the ground state has a net contribution to the charge due to $\phi_{0}$. Indeed, the definition of the electromagnetic current,
$j^{(em)}_{\mu}=-\frac{\partial L_{\phi A}}{\partial A^{\mu}}\,,$ (3.14)
gives
$j^{(em)}_{\mu}=iqJ_{\mu}\,,$ (3.15)
where
$J_{\mu}=\phi^{\ast}\hat{D}_{\mu}\phi-(\hat{D}_{\mu}\phi)^{\ast}\phi\,,$ (3.16)
and it is straightforward to verify that $j^{(em)}_{\mu}$ is conserved, as it should be. It then follows that the symmetry breaking mechanism produces a net current density given by
$j^{(em)}_{\mu}=q\mu\phi^{2}_{0}u_{\mu}\,.$ (3.17)
Physically, the total charge of the system must be zero, which is assured by assuming that the $\phi$ background is superimposed on an oppositely charged background to render the total system electrically neutral. Thus for practical purposes we can disregard that tadpole term.
4 Eigenmodes and the dispersion relations
4.1 The evolution equations
We write it in momentum space by making the correspondence[24],
$-\frac{1}{4}F^{2}$$\rightarrow$ $-V^{\ast\mu}k^{2}\tilde{g}_{\mu\nu}V^{\nu}\,,$
$\frac{1}{2}(\partial\rho)^{2}$$\rightarrow$ $k^{2}\rho^{\ast}\rho\,,$
$\frac{1}{2}m^{2}_{V}V^{2}$$\rightarrow$ $m^{2}_{V}V^{\ast\mu}V_{\mu}\,,$
$\frac{1}{2}m^{2}_{1}\rho^{2}$$\rightarrow$ $m^{2}_{1}\rho^{\ast}\rho\,,$
$(u\cdot V)\rho$$\rightarrow$ $(u\cdot V)^{\ast}\rho+c.c\,,$ (4.1)
and therefore
$L^{(2)}(k)=-V^{\ast\mu}[k^{2}\tilde{g}_{\mu\nu}-m^{2}_{V}g_{\mu\nu}]V^{\nu}+\rho^{\ast}(k^{2}-m^{2}_{1})\rho-2\mu m_{V}[(u\cdot V)^{\ast}\rho+c.c]\,,$ (4.2)
where
$\tilde{g}_{\mu\nu}=g_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{k^{2}}\,.$ (4.3)
If it were not for the last term, this is the usual free Lagrangian density for a vector with squared mass $m^{2}_{V}$ and a scalar with squared mass $m^{2}_{1}$. However the last term mixes the scalar with the longitudinal components of $V^{\mu}$.
We decompose $V^{\mu}$ in the form
$V^{\mu}=V^{\mu}_{T}+V_{L}e^{\mu}_{3}+\frac{k^{\mu}}{\sqrt{k^{2}}}V_{k}\,,\qquad V^{\mu}_{T}=\sum_{i=1,2}V_{i}e^{\mu}_{i}$ (4.4)
where
$e^{\mu}_{1,2}$ $=$ $(0,\vec{e}_{1,2})\,,$
$e^{\mu}_{3}$ $=$ $\frac{\tilde{u}^{\mu}}{\sqrt{-\tilde{u}^{2}}}\,,$ (4.5)
with
$\tilde{u}_{\mu}=\tilde{g}_{\mu\nu}u^{\nu}$ $=$ $u_{\mu}-\frac{(k\cdot u)k_{\mu}}{k^{2}}\,,$
$\sqrt{-\tilde{u}^{2}}$ $=$ $\frac{\kappa}{\sqrt{k^{2}}}\,.$ (4.6)
In particular,
$\tilde{u}\cdot V$ $=$ $-\sqrt{-\tilde{u}^{2}}V_{L}\,,$
$k\cdot V$ $=$ $\sqrt{k^{2}}V_{k}\,,$ (4.7)
and therefore
$u\cdot V=\tilde{u}\cdot V+\frac{(k\cdot u)(k\cdot V)}{k^{2}}=-\sqrt{-\tilde{u}^{2}}V_{L}+\frac{(k\cdot u)V_{k}}{\sqrt{k^{2}}}\,.$ (4.8)
Thus the quadratic part of the Lagrangian density is
$L^{(2)}(k)$ $=$ $-(k^{2}-m^{2}_{V})V^{\ast}_{T}\cdot V_{T}+(k^{2}-m^{2}_{V})V^{\ast}_{L}V_{L}+m^{2}_{V}V^{\ast}_{k}V_{k}+(k^{2}-m^{2}_{1})\rho^{\ast}\rho$ (4.9)
$\mbox{}+2\mu m_{V}\sqrt{-\tilde{u}^{2}}\left[V^{\ast}_{L}\rho+c.c.\right]-\frac{2\mu m_{V}(k\cdot u)}{\sqrt{k^{2}}}\left[V^{\ast}_{k}\rho+c.c\right]\,.$
Since the component $V_{k}$ does not have a kinetic energy term, and therefore it is not a dynamical variable, one way to proceed is to eliminate it from the Lagrangian by using its Lagrange equation. The equation is
$m^{2}_{V}V_{k}-\frac{2\mu m_{V}(k\cdot u)}{\sqrt{k^{2}}}\rho=0\,,$ (4.10)
and therefore
$V_{k}=\frac{2\mu(k\cdot u)}{m_{V}\sqrt{k^{2}}}\rho\,.$ (4.11)
Substituting this back in $L^{(2)}(k)$ we then get
$L^{(2)}(k)$ $=$ $-(k^{2}-m^{2}_{V})V^{\ast}_{T}\cdot V_{T}+(k^{2}-m^{2}_{V})V^{\ast}_{L}V_{L}+\left(k^{2}-m^{2}_{1}-\frac{4\mu^{2}(k\cdot u)^{2}}{k^{2}}\right)\rho^{\ast}\rho$ (4.12)
$\mbox{}+2\mu m_{V}\sqrt{-\tilde{u}^{2}}\left[V^{\ast}_{L}\rho+c.c.\right]\,.$
The alternative way is to find the equations for $V_{L}$ and $\rho$ and eliminate $V_{k}$ using Eq. (4.11). The final equations for $V_{L}$ and $\rho$ should be the same. As a verification, the Lagrange equations from Eq. (4.9) are,
$(k^{2}-m^{2}_{V})V_{L}+2\mu m_{V}\sqrt{-\tilde{u}^{2}}\rho$ $=$ $0\,,$
$(k^{2}-m^{2}_{1})\rho+2\mu m_{V}\sqrt{-\tilde{u}^{2}}V_{L}-2\mu m_{V}\frac{(k\cdot u)}{\sqrt{k^{2}}}V_{L}$ $=$ $0\,,$ (4.13)
and using Eq. (4.11) they become
$(k^{2}-m^{2}_{V})V_{L}+2\mu m_{V}\sqrt{-\tilde{u}^{2}}\rho$ $=$ $0\,,$
$2\mu m_{V}\sqrt{-\tilde{u}^{2}}V_{L}+\left[k^{2}-m^{2}_{1}-\frac{4\mu^{2}(k\cdot u)^{2}}{k^{2}}\right]\rho$ $=$ $0\,,$ (4.14)
which is the same that is obtained from Eq. (4.12). The picture that emerges is that the transverse modes develop a standard degenerate dispersion relation for the two polarizations, while the longitudinal mode $V_{L}$ mixes with the scalar $\rho$. Our next task is to solve the evolution equations we have obtained, to find the dispersion relations and the corresponding mixing parameters $V_{L}$ and $\rho$ of the eigenmodes.
4.2 Dispersion relations
Using Eq. (4.1), it follows that at $\kappa=0$ the equations in Eq. (4.1) decouple,
$(\omega^{2}-m^{2}_{V})V_{L}$ $=$ $0\,,$
$(\omega^{2}-m^{2}_{\rho})\rho$ $=$ $0\,,$ (4.15)
where
$m^{2}_{\rho}=m^{2}_{1}+4\mu^{2}\,.$ (4.16)
Thus we can identify the plasma frequency as
$\omega^{2}_{p0}=m^{2}_{V}=q^{2}\phi^{2}_{0}=\frac{q^{2}(\mu^{2}-m^{2})}{\lambda}\,,$ (4.17)
while the $\rho$ mode has mass $m^{2}_{\rho}$, which coincides with the mass of the “$+$” mode in the un-gauged model (see Eq. (2.26)).
For $\kappa\not=0$ the dispersion relations are the solutions of
$(k^{2}-m^{2}_{V})\left[k^{2}-m^{2}_{1}-\frac{4\mu^{2}(k\cdot u)^{2}}{k^{2}}\right]-\frac{4\mu^{2}m^{2}_{V}\kappa^{2}}{k^{2}}=0\,,$ (4.18)
which using Eq. (2) can be written as
$(k^{2}-m^{2}_{\rho})(k^{2}-m^{2}_{V})-4\mu^{2}\kappa^{2}=0\,.$ (4.19)
This is a quadratic equation for $k^{2}$, with solutions
$\omega^{2}=\kappa^{2}+\frac{1}{2}(m^{2}_{\rho}+m^{2}_{V})\pm\left[\frac{1}{4}(m^{2}_{\rho}-m^{2}_{V})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,.$ (4.20)
5 Discussion
For the following discussion it is convenient to normalize the various parameters as follows,
$x^{2}$ $=$ $\frac{\kappa^{2}}{2\mu^{2}}\,,$
$y^{2}$ $=$ $\frac{\omega^{2}}{2\mu^{2}}\,,$
$M^{2}$ $=$ $\frac{m^{2}_{\rho}+m^{2}_{V}}{4\mu^{2}}\,,$
$\Delta$ $=$ $\frac{|m^{2}_{\rho}-m^{2}_{V}|}{4\mu^{2}}\,.$ (5.1)
In terms of them, the dispersion relations given in Eq. (4.20) are expressed in the form
$y^{2}_{\pm}=x^{2}+M^{2}\pm\sqrt{\Delta^{2}+2x^{2}}\,.$ (5.2)
For guidance, plots of the dispersion relations are shown in Fig. 1 for sample values of the mass parameters. It must be kept in mind that numerically $\Delta<M^{2}$, and also it is important to recognize that
$M^{2}>1\,.$ (5.3)
Eq. (5.3) follows from the formulas for $m^{2}_{V}$ and $m^{2}_{\rho}$ in terms of $\mu$ (e.g. Eqs. (4.17) and (4.16)), that give
$M^{2}=1+\frac{1}{4\mu^{2}}\left[\frac{q^{2}(\mu^{2}-m^{2})}{\lambda}+2(\mu^{2}-m^{2})\right]\,.$ (5.4)
Remembering Eq. (2.8), this implies Eq. (5.3).
Figure 1: Sample plots of the dispersion relations as written in Eq. (5.2). We have set $M=2$, and $\Delta$ as indicated. The upper branches correspond to $\omega_{+}$ and the lower ones to $\omega_{-}$. The lowest branch, corresponding to $\omega_{-}$ with $\Delta=1$, is well approximated by a formula of the form given in Eq. (5.15) for small values of $\kappa$. See the discussion in the text.
From Eq. (5.2) it follows that $y_{+}$ increases steadily as $x$ increases, and $y_{+}>x$, for all values of $x$. In other words $\omega_{+}$ increases steadily with $\kappa$, and $\omega_{+}>\kappa$, for all values of $\kappa$. However, $\omega_{-}$ has some distinguishing features, and its behavior is different depending on the range of the value of $\kappa$. The plots in Figs. 2 and 3 show the dispersion relations covering different regimes of $\kappa$ to emphasize some of these features.
Firstly, it follows from Eq. (5.2) that
$y_{-}<x$ (5.5)
for
$x^{2}>\frac{1}{2}(M^{4}-\Delta^{2})\,.$ (5.6)
That is,
$\omega_{-}<\kappa\,,$ (5.7)
for
$\frac{\kappa^{2}}{\mu^{2}}>M^{4}-\Delta^{2}\,,$ (5.8)
as illustrated in Fig. 2. The dispersion relation in the $\kappa$ regime indicated in Eq. (5.8) is reminiscent of the photon dispersion relation in a medium with an index of refraction greater than 1.
Figure 2: Sample plots of the dispersion relations to illustrate that $\omega_{-}<\kappa$ when $\kappa$ is sufficiently large; i.e., Eqs. (5.6) and (5.8). For reference the line $y=x$ is shown as well. We have set $M=2$, and $\Delta$ as indicated.
Figure 3: Sample plots of the dispersion relation for $\omega_{-}$ illustrating the cases in which it has a miminum for certain range of values of the mass parameters. We have set $M=2$, and $\Delta$ as indicated.
Secondly, another noteworthy feature is the fact that if $\Delta<1$, then $\omega_{-}$ decreases in a certain range of $\kappa$, and reaching a minimum value. This is illustrated in Fig. 1 in the plots of $\omega_{-}$ for $\Delta=0.2,0.5$, and more emphatically in Fig. 3. In any case, the expression for $\omega^{2}_{-}$, as given by Eq. (5.2), remains positive for all values of $\kappa$. To show this let us note the following. From Eq. (5.2) it follows that $y^{2}_{-}$ becomes zero at $x=x_{\pm}$, where
$x^{2}_{\pm}=(1-M^{2})\pm\sqrt{(1-M^{2})^{2}-(M^{4}-\Delta^{2})}\,.$ (5.9)
Since $M^{2}>1$, as shown in Eq. (5.3), there is no solution of the equation
$y^{2}_{-}=0\,,$ (5.10)
and $\omega^{2}_{-}$ remains positive for all values of $\kappa$.
A related feature, in the case that $\Delta<1$, is that the group velocity of the $\omega_{-}$ branch becomes negative in the range of $\kappa$ in which $\omega_{-}$ decreases.
Figure 4: Sample plots of the group velocity $v_{g+}$ for the $\omega_{+}$ branch as written in Eq. (5.11). We have set $M=2$, and $\Delta$ as indicated.
Figure 5: Sample plots of the group velocity $v_{g-}$ for the $\omega_{-}$ branch as written in Eq. (5.11). We have set $M=2$, and $\Delta$ as indicated.
The group velocity for each branch is
$v_{g}=\frac{\partial\omega}{\partial\kappa}=\frac{x}{y}\left[1\pm\frac{1}{\sqrt{\Delta^{2}+2x^{2}}}\right]\,.$ (5.11)
Plots of $v_{g}$ are shown in Figs. 4 and 5 for the $\omega_{\pm}$ branches, respectively. Each figure also displays the corresponding plot of $v_{\pm}\equiv\kappa/\omega_{\pm}$ for reference. In each branch, $v$ does not change appreciably for the different values of $\Delta$. The group velocity of the $\omega_{-}$ branch becomes negative in the range of $\kappa$ in which $\omega_{-}$ decreases. In fact it is zero at
$x=\sqrt{\frac{1-\Delta^{2}}{2}}\,,$ (5.12)
and it is negative (positive) for $\kappa$ smaller (larger) than this value, respectively.
One particular feature worth mentioning is the result for the non-relativistic (NR) limit of the $\omega_{-}$ branch. For sufficiently small values of $\kappa$, the dispersion relations are
$(y^{2}_{\pm})_{NR}=(M^{2}\pm\Delta)+\left(1\pm\Delta^{-1}\right)x^{2}\mp\frac{x^{4}}{2\Delta^{3}}\,.$ (5.13)
Thus, in particular,
$(y^{2}_{-})_{NR}=(M^{2}-\Delta)+\left(1-\Delta^{-1}\right)x^{2}+\frac{x^{4}}{2\Delta^{3}}\,.$ (5.14)
Moreover, in the case that $\Delta\simeq 1$, the dispersion of the ()(-)( - ) branch is
$(y^{2}_{-})_{NR}\simeq(M^{2}-1)+\frac{1}{2}x^{4}\,.$ (5.15)
To state this more precisely, recall that in the NR limit, the chemical potential, which we denote by $\mu_{nr}$, is related to the chemical potential that we use in TFT by
$\mu=m+\mu_{nr}\,.$ (5.16)
This follows from the usual form of the distribution functions,
$f(p)=\frac{1}{e^{\beta(E(p)-\mu)}\pm 1}\,,$ (5.17)
and from the NR counterpart
$f_{NR}(p)=\frac{1}{e^{\beta(p^{2}/2m-\mu_{nr})}\pm 1}\,.$ (5.18)
Let us consider specifically the case
$m\gg\kappa,\mu_{nr},m_{V}\,.$ (5.19)
In that limit Eq. (5.14) gives
$(y^{2}_{-})_{NR}=\frac{m^{2}_{V}}{2\mu^{2}}+x^{2}O\left(\frac{\mu_{nr}}{m},\frac{m^{2}_{V}}{m^{2}}\right)+\frac{1}{2}x^{4}\left(1+O\left(\frac{\mu_{nr}}{m},\frac{m^{2}_{V}}{m^{2}}\right)\right)\,,$ (5.20)
or, in terms of the physical variables,
$(\omega^{2}_{-})_{NR}=m^{2}_{V}+\left(\frac{\kappa^{2}}{2m}\right)^{2}+O\left(\frac{\kappa^{2}\mu_{nr}}{m},\frac{\kappa^{2}m^{2}_{V}}{m^{2}}\right)\,.$ (5.21)
Thus if the $O\left(\kappa^{2}\mu_{nr}/m,\kappa^{2}m^{2}_{V}/m^{2}\right)$ terms can be neglected, the $\kappa^{2}$ term does not appear, which is reminiscent of the dispersion relation that has been obtained in NR models of the BE condensation of charged scalars[21]. However, in our case, the (±)plus-or-minus($\pm$)( ± ) dispersion relations are different, and this occurs only for the $\omega_{-}$ dispersion relation. More generally, the presence of a $O(\kappa^{2})$ term (even in the $T=0$ regime that we are considering) is mentioned in Ref. [22], as well as the fact that the coefficient is negative; which in our case can occur for $\omega_{-}$ if $\Delta<1$.
6 Polarization tensor and dielectric constant
An alternative way to obtain the dispersion relations is through the photon polarization tensor, or self-energy. This method has also the virtue that it provides us with an expression for the dielectric function of the system, which is useful on its own merits. The starting point is the bilinear part of the Lagrangian density for $V_{\mu}$ and $\rho$ given in Eq. (4.12). The idea is to integrate out the variable $\rho$, which is the tree-level analog of calculating the 1-loop self-energy, leaving only the dynamical variables of $V$, and identify the polarization tensor by comparing the result with
$L_{eff}=-V^{\ast\mu}\left[(k^{2}-\pi_{T})R_{\mu\nu}+(k^{2}-\pi_{L})Q_{\mu\nu}\right]V^{\nu}\,,$ (6.1)
where $R_{\mu\nu}$ and $Q_{\mu\nu}$ are defined by
$Q_{\mu\nu}$ $=$ $\frac{\tilde{u}_{\mu}\tilde{u}_{\nu}}{\tilde{u}^{2}}\,,$
$R_{\mu\nu}$ $=$ $\tilde{g}_{\mu\nu}-Q_{\mu\nu}\,.$ (6.2)
In terms of the variables $V_{T,L}$ we have defined in Eq. (4.4),
$L_{eff}=-(k^{2}-\pi_{T})V^{\ast}_{T}\cdot V_{T}+(k^{2}-\pi_{L})V^{\ast}_{L}V_{L}\,.$ (6.3)
From Eq. (4.12), the equation for $\rho$ is precisely the equation given in the second line of Eq. (4.1). Solving it for $\rho$ and substituting the solution back in Eq. (4.12) gives
$L^{(2)}(k)=-(k^{2}-m^{2}_{V})V^{\ast}_{T}\cdot V_{T}+(k^{2}-m^{2}_{V})V^{\ast}_{L}V_{L}-\frac{(2\mu m_{V}\sqrt{-\tilde{u}^{2}})^{2}}{k^{2}-m^{2}_{1}-\frac{4\mu^{2}(k\cdot u)^{2}}{k^{2}}}V^{\ast}_{L}V_{L}\,.$ (6.4)
Comparing Eqs. (6.3) and (6.4), we then identify
$\pi_{L}=m^{2}_{V}+\frac{(2\mu m_{V}\sqrt{-\tilde{u}^{2}})^{2}}{k^{2}-m^{2}_{1}-\frac{4\mu^{2}(k\cdot u)^{2}}{k^{2}}}\,,$ (6.5)
which using Eqs. (2), (4.1) and (4.16) can be written in the form
$\pi_{L}=m^{2}_{V}+\frac{4\mu^{2}m^{2}_{V}\kappa^{2}}{k^{2}(k^{2}-m^{2}_{\rho})-4\mu^{2}\kappa^{2}}\,.$ (6.6)
It is simple to verify that the dispersion relation equation
$k^{2}-\pi_{L}=0\,,$ (6.7)
gives again Eq. (4.18), as it should be. While this is reassuring, the virtue of Eq. (6.6) is that it gives us an expression for the longitudinal dielectric function, as function of $\omega$ and $\kappa$,
$\epsilon_{\ell}=1-\frac{\pi_{L}}{k^{2}}\,,$ (6.8)
which allows to study also the response of the system to external fields.
In this context we would like to emphasize the following. What we have done is to find the modes that diagonalize the scalar-photon part of the Lagrangian, in the BE background. The dispersion relations that we have written in Eq. (4.20) are the dispersion relations of those modes.
In reality, as we have mentioned, the BE condensate must be superimposed on an oppositely charged background component. It would be a multi-component system, of which one of them is the BE condensate. Both components contribute to the total $\pi_{L}$ (as well as the dielectric constant), and therefore the dispersion relations of the propagating plasma modes have contributions from both components. Physically the propagating modes are combinations of the BE modes and the plasma modes of the background component.
In systems in which the dominant contribution to $\pi_{L}$ arises from the BE component [which we have written in Eq. (6.6)], then the dispersion relation of the propagating photon (obtained by solving Eq. (6.7)) is given by the dispersion relation of the BE modes that we have written in Eq. (4.20).
In more general cases, $\pi_{L}$ would have an additional contribution besides the BE contribution of Eq. (6.6), depending on the composition and conditions of the background component, and the dispersion relations of the propagating modes would be correspondingly different from those given in Eq. (4.20).
The main result is that, in such a multicomponent system, Eq. (6.6) represents the contribution to $\pi_{L}$ (and whence to the dielectric constant) due to the BE component, which can be used to study the implications for the optical properties of the system, which are different from what they would be if the system does not have the BE component.
7 Mixing
Besides the dispersion relation and the polarization tensor, another important factor is the $\rho-V_{L}$ mixing in the eigenmodes. Once the solutions $\omega_{\pm}$ are given, these are determined from Eq. (4.1). From Eq. (4.20) for the dispersion relations it follows that for $\kappa=0$,
$\omega^{2}_{+}(0)$ $=$ $m^{2}_{>}\,,$
$\omega^{2}_{-}(0)$ $=$ $m^{2}_{<}\,,$ (7.1)
where $m_{>}$ ($m_{<}$) is the greater (lesser) of $m_{\rho}$ and $m_{V}$. In what follows we consider the case $m_{\rho}>m_{V}$. The opposite case can be treated similarly.
Figure 6: Plots of the mixing coefficients $\rho_{+},V_{L+}$ defined in Eq. (7.6), as functions of $\kappa$. We have set $M=2$, and $\Delta=1$.
Figure 7: Plots of the mixing coefficients $\rho_{-},V_{L-}$ defined in Eq. (7.6) as functions of $\kappa$. We have set $M=2$, and $\Delta=1$.
A convenient way to write the elements of the mode eigenvectors in this case is
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}$ $=$ $\frac{1}{\sqrt{N_{-}}}\left(\begin{array}[]{c}m_{V}\gamma_{-}\\[12.0pt] \Lambda+\gamma^{2}_{-}\end{array}\right)\,,$ (7.6)
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{+}$ $=$ $\frac{1}{\sqrt{N_{+}}}\left(\begin{array}[]{c}\Lambda\\[12.0pt] -m_{V}\gamma_{+}\end{array}\right)\,,$ (7.11)
where
$\gamma_{\pm}$ $=$ $\frac{2\mu\kappa}{\sqrt{k^{2}_{\pm}}}\,,$
$k^{2}_{\pm}$ $=$ $\omega^{2}_{\pm}-\kappa^{2}=\frac{1}{2}(m^{2}_{\rho}+m^{2}_{V})\pm\left[\frac{1}{4}(m^{2}_{\rho}-m^{2}_{V})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,,$
$\Lambda$ $=$ $\frac{1}{2}(m^{2}_{\rho}-m^{2}_{V})+\left[\frac{1}{4}(m^{2}_{\rho}-m^{2}_{V})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,,$ (7.12)
and
$N_{-}$ $=$ $(\Lambda+\gamma^{2}_{-})^{2}+m^{2}_{V}|\gamma_{-}|^{2}\,,$
$N_{+}$ $=$ $\Lambda^{2}+m^{2}_{V}\gamma^{2}_{+}\,.$ (7.13)
With this convention, in the limit $\kappa=0$ these solutions reduce to
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}$ $=$ $\left(\begin{array}[]{c}0\\ 1\end{array}\right)\,,$ (7.18)
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{+}$ $=$ $\left(\begin{array}[]{c}1\\ 0\end{array}\right)\,,$ (7.23)
when $m_{\rho}>m_{V}$. Plots of the mixing coefficients $|\rho_{\pm}|^{2}$ and $|V_{L\pm}|^{2}$ is shown in Fig. 6 and Fig. 7. For the plots we take $M^{2}$ and $\Delta$ as indicated in the caption, and we consider the case $m_{\rho}>m_{V}$, so that, in particular, $(m_{V}/2\mu)^{2}=\frac{1}{2}(M^{2}-\Delta)$. The plots do not vary significantly for other values of $M^{2}$ and $\Delta$.
As already indicated above, at $\kappa=0$ the $\omega_{-}$ solution corresponds to a pure $V_{L}$,
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\left(\begin{array}[]{c}0\\ 1\end{array}\right)\qquad(\kappa=0)\,.$ (7.24)
At very large values of $\kappa$, it follows from Eq. (7.6) that the $\rho$ component dominates in both branches, which is reflected in the plots in Figs. 6 and 7.
On the other hand, the dispersion relation $\omega_{-}$ satisfies
$k^{2}_{-}=0\,,$ (7.25)
for $\kappa=\kappa_{0}$, where
$\kappa_{0}=\frac{m_{V}m_{\rho}}{2\mu}\,.$ (7.26)
It is easy to see from Eq. (7), that $\gamma_{-}$ increases relative to $\Lambda$ as $k^{2}_{-}\rightarrow 0$, and therefore from Eq. (7.6) it follows that for $\kappa=\kappa_{0}$ the ()(-)( - ) mode is a pure $V_{L}$
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\left(\begin{array}[]{c}0\\ 1\end{array}\right)\qquad(\kappa=\kappa_{0})\,.$ (7.27)
Moreover, $k^{2}_{-}$ is positive for $\kappa<\kappa_{0}$, but it is negative for $\kappa>\kappa_{0}$. Therefore, for $\kappa>\kappa_{0}$, we have
$\gamma_{-}=-\frac{2i\mu\kappa}{\sqrt{|k^{2}_{-}|}}\,,$ (7.28)
and then the $\omega_{-}$ eigenspinor is of the form
$\left(\begin{array}[]{c}\rho\\ V_{L}\end{array}\right)_{-}=\frac{1}{\sqrt{N_{-}}}\left(\begin{array}[]{c}-im_{V}|\gamma_{-}|\\[12.0pt] \Lambda-|\gamma_{-}|^{2}\,.\end{array}\right)\qquad(\kappa>\kappa_{0})\,.$ (7.29)
The situation here resembles the circular polarization of a propagating photon, or more precisely in this case, elliptic polarization,
$e_{\pm}=\left(\begin{array}[]{c}\pm ia\\ b\end{array}\right)\,.$ (7.30)
Thus in the present case, if $\kappa>\kappa_{0}$, the $\omega_{-}$ mode oscillates between the two flavors ($V_{L}$ and $\rho$), as it propagates, and at some spacetime points it is a pure $V_{L}$ and at other points a pure $\rho$.
8 Conclusions and Outlook
We have presented a model for the propagation of photons in a medium that contains a Bose-Einstein (BE) condensate of a charged scalar field. The idea is based on the model recently proposed to treat the propagation of fermions in a BE condensate, extended by coupling the scalar field with the electromagnetic field in the usual way. Our goal has been to study the effects of the symmetry breaking on the photon dispersion relations and the optical properties of the system. To this end, the propagating modes were identified, and the corresponding dispersion relations were determined, in Section 4. As we have shown there, two of the modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. There are two other modes, to which we refer as the (±)plus-or-minus($\pm$)( ± ) modes, that are combinations of the longitudinal photon and the massive scalar field. Some salient aspects of the results obtained for the dispersion relations of the (±)plus-or-minus($\pm$)( ± ) modes were discussed in Section 5. The dispersion relation of the (+)(+)( + ) mode is always greater than the momentum, and increases steadily as the momentum increases. The behavior of the dispersion relation for the ()(-)( - ) mode is very different. On one hand, it decreases as the momentum increases in a given momentum range, and therefore the corresponding group velocity is negative in that range. In another range, for values of momentum above a given threshold, it increases steadily, but such that it remains smaller than the momentum as it increases, akin to a medium with index of refraction greater than 1. In Section 5 we also considered the non-relativistic limit of the dispersion relations, and discussed some aspects of the results in particular for the ()(-)( - ) mode. As shown there, under some conditions the dispersion relation of the ()(-)( - ) mode reduces to the form of the dispersion relation that has been obtained in non-relativistic models of the BE condensation of charged scalars[21, 22]. As a complement, in Section 6 we gave the expression for the contribution to the photon polarization tensor due to the BE condensate. This provides an alternative way to obtain the dispersion relations, but more importantly it has the virtue that it gives the corresponding contribution to the dielectric function of the system, which is useful on its own merits, e.g. for studying also the response of the system to external fields. Another useful result is the determination of the mixing parameters of the (±)plus-or-minus($\pm$)( ± ) modes presented in Section 7, which can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system.
In summary, altogether, the results we have obtained show that the BE condensate model we have considered has some particular non-trivial optical properties. Our aim has been to analyze them, in the context of the model, with the expectation that the model and the results obtained can be useful in some of the specific application contexts considered in the references cited.
The work of S. S. is partially supported by DGAPA-UNAM (Mexico) PAPIIT project No. IN103522.
Appendix A Justification of Eq. (4.1)
The justification of Eq. (4.1) is the following. For any field we write
$a(x)=\int\frac{d^{4}k}{(2\pi)^{4}}a(k)e^{-ik\cdot x}\,.$ (A.1)
If $a(x)$ is real (Hermitian), then
$a(-k)=a^{\ast}(k)\,.$ (A.2)
Now consider something like (an action) involving two real fields
$S=\int d^{4}x\left[\frac{1}{2}a^{2}(x)+\frac{1}{2}b^{2}(x)+\alpha b(x)a(x)\right]\,.$ (A.3)
What is the Lagrangian density in momentum space? Using the transformation of $a$ and $b$,
$S$ $=$ $\int\frac{d^{4}k}{(2\pi)^{4}}\int\frac{d^{4}k^{\prime}}{(2\pi)^{4}}\left[\frac{1}{2}a(k^{\prime})a(k)+\frac{1}{2}b(k^{\prime})b(k)+\alpha b(k^{\prime})a(k)\right]\delta(k+k^{\prime})$ (A.4)
$=$ $\int\frac{d^{4}k}{(2\pi)^{4}}\left[\frac{1}{2}a(-k)a(k)+\frac{1}{2}b(-k)b(k)+\alpha b(-k)a(k)\right]\,.$
But we would not obtain the correct equations if we just use
$L(k)=\frac{1}{2}a(-k)a(k)+\frac{1}{2}b(-k)b(k)+\alpha b(-k)a(k)\,,$ (A.5)
which is equal to
$L(k)=\frac{1}{2}a^{\ast}(k)a(k)+\frac{1}{2}b^{\ast}(k)b(k)+\beta b^{\ast}(k)a(k)\,.$ (A.6)
In particular, notice that this $L(k)$ is not real (Hermitian). The correct procedure is to notice that, since the first term mixes $a(k)$ with $a(-k)$, and similarly with $b(k)$, we must include both also in the $\alpha$ term. Therefore we must write
$S=\int_{k>0}\frac{d^{4}k}{(2\pi)^{4}}\left[\frac{1}{2}a(-k)a(k)+\frac{1}{2}b(-k)b(k)+\alpha b(-k)a(k)+(k\leftrightarrow-k)\right]\,,$ (A.7)
and therefore the correct Lagrangian density to use in momentum space is
$L(k)=a^{\ast}(k)a(k)+b^{\ast}(k)b(k)+\alpha\left(b^{\ast}(k)a(k)+c.c\right)\,,$ (A.8)
with the understanding that the complete action is
$S=\int_{k>0}\frac{d^{4}k}{(2\pi)^{4}}L(k)\,.$ (A.9)